OpenStudy (anonymous):

how to solve for x if the equetion is 3(x-2)squre =9

OpenStudy (smurfy14):

first divide by 3 (x-2)^2=9

OpenStudy (anonymous):

*=3 RHS ?

OpenStudy (smurfy14):

do (x-2)(x-2) and you get x^2-4x+4=9 then subtract 9 to make it x^2-4x-5

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

nope the answer is

OpenStudy (smurfy14):

oh just kidding!

OpenStudy (anonymous):

(x-2)^2 = 3 x-2 = +/- root 3 so x = 2 +/- root 3

OpenStudy (anonymous):

\[x=2\pm \sqrt{3}\]

OpenStudy (smurfy14):

ya i kinda messed up sorry :/

OpenStudy (anonymous):

that's okay